OpenStudy (anonymous):

In this lab there is a video about 2 seconds long. A tennis ball bounces in the screen and then dissapears again in about 1.5 seconds. I need to find the initial velocity. There is x,y graph taped to a wall with the black tape marking where the black lines are located. The distance between the two vertical black lines in the video is 50 cm. The distance between the two horizontal black lines in the video is 50 cm; but do not assume the same scale for the horizontal as for the vertical. The frame rate of the video is 30 frames per second. Assume that position (0, 0) is the bottom, left corner.

OpenStudy (anonymous):

so i know i have to use these equations..... v^2=v(i)^2+2ax v(i)=initial velocity a=-9.8 how would i find the initial velocity???????? sooo confused!

OpenStudy (abmon98):

you can use v+u/t=s